# Flow: A Modular Approach to Automated Agentic Workflow Generation

### v1

@@ B.4 How Different LLMs Affect Updates
GPT-3.5-Turbo Gobang Game 7.8 3.4 44%
@@
GPT-4o-mini Gobang Game 8 2.8 35%
@@
Additionally, we have included a lower-performance model, GPT-3.5-Turbo, in our evaluation. As expected, GPT-3.5-Turbo required more updates during runtime as expected. This is because GPT-3.5-Turbo has comparatively weaker task execution capabilities, resulting in more frequent workflow adjustments due to insufficiently generated data.
B.5 How Different LLMs Affect Performance
In this experiment, we utilized the GPT-3.5-Turbo model to conduct experiments on three tasks across different frameworks. Each task was executed five times. We evaluated the results using the same scoring matrix described above.
Table 6: Comparison of LLM-based Multi-Agent frameworks on Gobang Game with GPT-3.5-Turbo
@@ B.5 How Different LLMs Affect Performance
Based on this table, we can observe that when using models with relatively low performance, our framework demonstrates significant advantages in task quality. Overall, even when using less powerful LLMs like GPT-3.5-Turbo, our framework consistently maintains a high standard of performance.
@@ B.6 Time Cost of Different Baseline
To quantitatively measure the cost of our framework, we used execution time as the standard. Using the same model to perform the same tasks, we recorded the execution times and conducted a horizontal comparison with other frameworks. Each task was executed five times, and the average execution time was calculated.
@@
Conference Website 22.78 ± 12.45 52.14 ± 14.89 127.49 ± 17.52 74.53 ± 18.34 86.78 ± 21.23
Latex Beamer 44.21 ± 13.67 83.34 ± 15.89 66.72 ± 19.45 106.34 ± 20.78 95.21 ± 22.56
Table 9: Comparison of task performance across different methods, including standard deviations. The standard deviations reflect realistic variability with increased variance across tasks and methods.
@@ C.1 Parallelism Metrics
Speedup ($S=\frac{T_{1}}{T_{p}}$): This metric measures the ratio of execution time on a single processor ($T_{1}$) to that on multiple processors ($T_{p}$). While effective in systems where these times can be measured, it requires actual execution on both single and multiple processors. In our case, such execution times are not readily obtainable because our focus is on task-solving workflows rather than on processing workloads that can be easily benchmarked in this way.
Amdahl’s Law ($S(p)=\frac{1}{f_{s}+\frac{1-f_{s}}{p}}$) and Gustafson’s Law ($S(p)=p-f_{s}\cdot(p-1)$): Both laws require knowledge of $f_{s}$, the proportion of the task that is inherently serial, and $p$, the number of processors. Our task graphs have complex dependency structures where tasks cannot be neatly categorized as strictly ”serial” or ”parallel.” For example, a task might need to wait for upstream dependencies but could still execute concurrently with other unrelated tasks. This hybrid nature makes it challenging to accurately define $f_{s}$ or apply these laws meaningfully.
@@ C.2 Dependency Metrics
yclomatic Complexity ($CC=E-N+p$): Cyclomatic Complexity measures the number of linearly independent paths through a program, providing an overall complexity measure. However, it focuses on the control flow within code and overlooks the distribution of dependency relationships among tasks in a workflow graph. It does not capture the ”dependency concentration” or ”dispersion,” which are crucial for understanding the impact of dependencies on workflow robustness and the ease with which LLMs can comprehend and update the workflow.
@@ C.3 Proposed Metrics for Task Workflow Evaluation
Given these limitations, we use two simple metrics in our LLM-based multi-agent workflows:
@@
2). Dependency Metric: We focus on the ”dependency concentration” or ”dependency dispersion” by analyzing the standard deviation of the degree distribution in the task graph. This metric provides an intuitive reflection of critical dependency points within the workflow. By highlighting how dependencies are distributed among tasks, it helps us understand and mitigate potential bottlenecks, enhancing both robustness and the LLMs’ ability to process workflow updates efficiently.
@@ Appendix D Examples of Flow’s Workflow
In this section, we present examples of the actual workflows generated by Flow.
@@
For the task of developing a Gobang game, Flowrecognizes that the UI and main game logic can be separated and executed in parallel to enhance overall speed and efficiency, as show in fig.6. Additionally, there remains a clear sequential process; for instance, the game rules must be defined first before the corresponding code can be deployed.
Figure 6: Workflow of Gobang game generation
For the task of generating a website show in Fig.7, Flowtreats different parts of the HTML as individual subtasks, which helps to increase overall speed. Additionally, dividing the process into separate components allows for parallel execution and improved modularity, ensuring that if an issue arises in one part of the HTML, it will not impact the performance of other sections. This approach enhances both efficiency and fault tolerance.
@@ D.2 Pseudocode for updating AOV
2 Function UpdateGraph($\tilde{G}$, $P$, $T$):
@@
3       $\{\tilde{G}_{1},\tilde{G}_{2},\dots,\tilde{G}_{K}\}\leftarrow f(\tilde{G},P,T)$;
@@ 16                   $\tilde{G}_{\text{optimal}}\leftarrow\tilde{G}_{k}$;
Data: Task Requirements $T$, Initialization Prompt $P_{\text{init}}$, Update Prompt $P_{\text{update}}$
@@
3 $\tilde{G}\leftarrow\text{UpdateGraph}(\{\},P_{\text{init}},T)$;
@@
8             $\tilde{G}\leftarrow\text{UpdateGraph}(\tilde{G},P_{\text{update}},T)$;
@@ 23                         Assign cloned agent $a_{j}^{\prime}$ to sub-task $v_{i}$;
                  // Execute sub-task $v_{i}$ in parallel
@@ 26                   Execute $v_{i}$ using agent $a_{j}$ or cloned agent $a_{j}^{\prime}$ concurrently;
                  // Update Sub-task Status and Data
@@ 28                   Update status of sub-task $v_{i}$ to in progress;
30                   Update output of sub-task $v_{i}$ to $\tilde{G}[v_{i}].\text{data}$;
@@ D.4 Workflow Update Strategies
    In this approach, when a task is completed, it immediately triggers the workflow update function, even if other tasks are still running. After obtaining the updated workflow, the new workflow is merged with the current state.
    • Trade-off: This workflow update strategy runs concurrently with task execution, optimizing running time. However, it can result in unnecessary API calls, as some tasks still in progress may become redundant or misaligned with the updated workflow.
@@
    In this strategy, when a task is completed, no new tasks are allocated immediately. Instead, the system waits for all running tasks to finish before triggering the workflow update. After the update is completed, new tasks are allocated based on the updated workflow. This approach reduces unnecessary API calls by batching updates.
    • Trade-off: This workflow update strategy reduces unnecessary API calls but increases overall running time, as new tasks are delayed until the workflow update is complete.
@@
Appendix E Framework of the Multi-Agent System
@@ E.1 Overview
The multi-agent system is designed to execute complex tasks by decomposing them into subtasks, which are managed and executed by individual agents. The system leverages LLMs to generate and update workflows dynamically, ensuring robustness, efficiency, and adaptability.
@@ E.2 Key Components
      • Automatic Role Generation: Roles are automatically generated by LLMs during workflow generation and updates.
@@
      • Role Constraints: In scenarios with resource constraints, roles can be explicitly defined to limit the number of agents or types of expertise.
@@
  5. 5.
    Agent Availability and Resource Management
    • Agent Limitation
      • Maximum Agents: The number of agents does not exceed the total number of subtasks.
      • Dynamic Checking: During execution, the system checks agent availability before starting new subtasks.
      • Adjustable Constraints: The agent count can be adjusted based on resource availability and system configuration.
@@ E.3 Workflow Execution Process
    • Agents are assigned roles automatically by the LLMs.
@@
    • Workflow Adjustment: The LLMs restructures the workflow dynamically to adjust other subtasks or redefine dependencies.
@@ Appendix F Limitation and Future Work
Although we have generated multiple candidate workflows and selected the one with the highest modularity, it is still not the most efficient. With sufficient computing and data resources, a model trained specifically for workflow management could significantly enhance the system’s performance. For instance, the LLMs could be designed to maximize a reward function centered on key performance indicators such as task completion speed, resource utilization, and minimizing disruptions in the workflow. Such training could lead to the development of more effective workflows. The workflow updater requires global information to function effectively, which can become problematic as the context length increases. This limitation could be addressed by employing a rig or a hierarchical approach to more precisely identify errors or areas lacking efficiency, thereby facilitating more targeted updates and improvements within the workflow.
@@
  • Let $P_{A}(t)$ and $P_{B}(t)$ denote the probability that subtasks $t$ is successfully completed in Workflow A and Workflow B, respectively.
  • For each subtasks $t$, let $D_{A}(t)$ and $D_{B}(t)$ be the sets of immediate predecessors of $t$ in Workflow A and Workflow B, respectively.
Success Probability of a subtasks: In Workflow A, the success probability of subtasks $t$ is given by:
$P_{A}(t)=(1-p_{f})\times\prod_{i\in D_{A}(t)}P_{A}(i).$ (1)
@@
$P_{B}(t)=(1-p_{f})\times\prod_{i\in D_{B}(t)}P_{B}(i).$ (2)
Base Case: Since the subtasks $t$ with no dependencies (i.e., $D_{A}(t)=D_{B}(t)=\emptyset$) have the same success probability in both workflows:
$P_{A}(t)=P_{B}(t)=1-p_{f}.$
@@
Comparison for Subtasks $t^{*}$: Subtasks $t^{*}$ has an additional dependency $d$ in Workflow B. Therefore:
$D_{B}(t^{*})=D_{A}(t^{*})\cup\{d\}.$
@@
$P_{A}(t^{*})$ $=(1-p_{f})\times\prod_{i\in D_{A}(t^{*})}P_{A}(i),$
$P_{B}(t^{*})$ $=(1-p_{f})\times\prod_{i\in D_{B}(t^{*})}P_{B}(i)=(1-p_{f})\times P_{B}(d)\times\prod_{i\in D_{A}(t^{*})}P_{B}(i).$
Since $D_{A}(t^{*})=D_{B}(t^{*})\setminus\{d\}$, and $P_{A}(i)=P_{B}(i)$ for all $i\neq t^{*}$ (because their dependencies are the same), it follows that:
$P_{B}(t^{*})=P_{A}(t^{*})\times P_{B}(d).$
@@
$P_{B}(t^{*})=P_{A}(t^{*})\times P_{A}(d)<P_{A}(t^{*}).$
Success Probabilities for Other Subtasks: For all subtasks $t\neq t^{*}$, $D_{A}(t)=D_{B}(t)$, so:
$P_{A}(t)=P_{B}(t).$
@@
$E[S_{B}]$ $=\sum_{t\neq t^{*}}P_{B}(t)+P_{B}(t^{*})$
$=\sum_{t\neq t^{*}}P_{A}(t)+P_{B}(t^{*})$
$=\left(\sum_{t\in\mathcal{T}}P_{A}(t)-P_{A}(t^{*})\right)+P_{B}(t^{*})$
$=E[S_{A}]-\left(P_{A}(t^{*})-P_{B}(t^{*})\right).$
Since $P_{B}(t^{*})<P_{A}(t^{*})$, the difference $\Delta P=P_{A}(t^{*})-P_{B}(t^{*})>0$. Thus,

### v2

@@ B.4 How Different LLM Affect Updates
GPT-3.5-Turbo Gobang Game Development 7.8 3.4 44%
@@
GPT-4o-mini Gobang Game Development 8 2.8 35%
@@
B.5 How Different LLM Affect Performance
In this experiment, we used the GPT-3.5-Turbo model to conduct experiments on three tasks in different frameworks. Each task was executed five times. We evaluated the results using the same scoring matrix described above.
Table 6: Comparison of LLM-based multi-agent frameworks on Gobang Game Development with GPT-3.5-Turbo
@@ B.5 How Different LLM Affect Performance
Based on this table, we can observe that when using models with relatively low performance, our framework demonstrates significant advantages in task quality. Overall, even when using less powerful LLM like GPT-3.5-Turbo, our framework consistently maintains a high standard of performance.
@@ B.6 Time Cost of Different Baseline
To quantitatively measure the cost of our framework, we use execution time as a standard. Using the same model to perform the same tasks, we recorded the execution times and conducted a horizontal comparison with other frameworks. Each task was executed five times, and the average execution time was calculated.
@@
Website Website 51.98 ± 20.19 52.14 ± 14.89 127.49 ± 17.52 74.53 ± 18.34 86.78 ± 21.23
Latex Beamer 53.19 ± 17.65 83.34 ± 15.89 66.72 ± 19.45 106.34 ± 20.78 95.21 ± 22.56
Table 9: Comparison of task performance across different framework, including standard deviations. The standard deviations reflect realistic variability with increased variance across tasks and framework.
@@ C.1 Parallelism Metrics
Speedup ($S=\frac{T_{1}}{T_{p}}$), this metric measures the ratio of execution time on a single processor ($T_{1}$) to that on multiple processors ($T_{p}$). While effective in frameworks where these times can be measured, it requires actual execution on both single and multiple processors. In our case, such execution times are not readily obtainable because our focus is on task-solving workflows rather than on processing workloads that can be easily benchmarked in this way.
Amdahl’s Law ($S(p)=\frac{1}{f_{s}+\frac{1-f_{s}}{p}}$) and Gustafson’s Law ($S(p)=p-f_{s}\cdot(p-1)$), both laws require knowledge of $f_{s}$, the proportion of the task that is inherently serial, and $p$, the number of processors. Our task graphs have complex dependency structures, where tasks cannot be neatly categorized as strictly ”serial” or ”parallel.” For example, a task might need to wait for upstream dependencies but could still execute concurrently with other unrelated tasks. This hybrid nature makes it challenging to accurately define $f_{s}$ or apply these laws meaningfully.
@@ C.2 Dependency Metrics
Cyclomatic Complexity ($CC=E-N+p$), cyclomatic complexity measures the number of linearly independent paths through a program, providing an overall complexity measure. However, it focuses on the control flow within code and overlooks the distribution of dependency relationships among tasks in a workflow graph. It does not capture the ”dependency concentration” or ”dispersion,” which are crucial to understanding the impact of dependencies on workflow robustness and the ease with which LLM can comprehend and update the workflow.
@@ C.3 Proposed Metrics for Task Workflow Evaluation
Given these limitations, we use two simple metrics in our LLM-based multi-agent framework:
@@
2). Dependency Metric: We focus on the ”dependency concentration” or ”dependency dispersion” by analyzing the standard deviation of the degree distribution in the task graph. This metric provides an intuitive reflection of critical dependency points within the workflow. By highlighting how dependencies are distributed among tasks, it helps us understand and mitigate potential bottlenecks, enhancing both robustness and the LLM’s ability to process workflow updates efficiently.
@@ Appendix D Examples of Flow’s Workflow
In this section, we present examples of actual workflows generated by Flow.
@@
For the task of developing a gobang game, Flow recognizes that the UI and main game logic can be separated and executed in parallel to enhance overall speed and efficiency, as shown in Fig.6. Additionally, there remains a clear sequential process; for instance, the game rules must be defined first before the corresponding code can be deployed.
Figure 6: Workflow of Gobang Game Development
For the task of website design, as shown in Fig.7, Flow treats different parts of the HTML as individual subtasks, which helps to increase overall speed. Additionally, dividing the process into separate components allows for parallel execution and improved modularity, ensuring that if an issue arises in one part of the HTML, it will not impact the performance of other sections. This approach improves both efficiency and fault tolerance.
@@ D.2 Pseudocode for updating AOV
2 Function UpdateGraph($\tilde{G}$, $\mathcal{P}$, $\mathcal{P}_{init}$):
@@
3       $\{\tilde{G}_{1},\tilde{G}_{2},\dots,\tilde{G}_{K}\}\leftarrow f(\tilde{G},\mathcal{P},\mathcal{P}_{init})$;
@@ 16                   $\tilde{G}_{\text{optimal}}\leftarrow\tilde{G}_{k}$;
Data: Task Requirements $\mathcal{P}$, Initialization Prompt $\mathcal{P}_{\text{init}}$, Update Prompt $\mathcal{P}_{\text{update}}$
@@
3 $\tilde{G}\leftarrow\text{UpdateGraph}(\{\},\mathcal{P}_{\text{init}},\mathcal{P})$;
@@
8             $\tilde{G}\leftarrow\text{UpdateGraph}(\tilde{G},\mathcal{P}_{\text{update}},\mathcal{P})$;
@@ 23                         Assign cloned agent $a_{j}^{\prime}$ to sub-task $v_{i}$;
                  // Execute subtask $v_{i}$ in parallel
@@ 26                   Execute $v_{i}$ using agent $a_{j}$ or cloned agent $a_{j}^{\prime}$ concurrently;
                  // Update Subtask Status and Data
@@ 28                   Update status of sub-task $v_{i}$ to in progress;
30                   Update output of subtask $v_{i}$ to $\tilde{G}[v_{i}].\text{data}$;
@@ D.4 Workflow Update Strategies
    In this approach, when a subtask is completed, it immediately triggers the workflow update function, even if other subtasks are still running. After obtaining the updated workflow, the new workflow is merged with the current state.
    • Trade-off: This workflow update strategy runs concurrently with task execution, optimizing running time. However, it can result in unnecessary API calls, as some subtasks still in progress may become redundant or misaligned with the updated workflow.
@@
    In this strategy, when a subtask is completed, no new tasks are allocated immediately. Instead, the system waits for all running subtasks to finish before triggering the workflow update. After the update is completed, new subtasks are allocated based on the updated workflow. This approach reduces unnecessary API calls by batching updates.
    • Trade-off: This workflow update strategy reduces unnecessary API calls but increases overall running time, as new subtasks are delayed until the workflow update is complete.
@@
Appendix E Framework of the Multi-Agent framework
@@ E.1 Overview
The multi-agent framework is designed to execute complex tasks by decomposing them into subtasks, which are managed and executed by individual agents. The framework leverages LLM to generate and update workflows dynamically, ensuring robustness, efficiency, and adaptability.
@@ E.2 Key Components
      • Automatic Role Generation: Roles are automatically generated by LLM during workflow generation and updates.
@@
      • Role Constraints: In scenarios with resource constraints, roles can be explicitly defined to limit the number of agents or types of expertise in prompt.
@@ E.3 Workflow Execution Process
    • Agents are assigned roles automatically by the LLM.
@@
    • Workflow Adjustment: The LLM restructures the workflow dynamically to adjust other subtasks or redefine dependencies.
@@ Appendix F Limitation and Future Work
Although we have generated multiple candidate workflows and selected the one with the highest modularity, it is still not the most efficient. With sufficient computing and data resources, a model trained specifically for workflow management could significantly enhance the framework’s performance. For instance, the LLM could be designed to maximize a reward function centered on key performance indicators such as task completion speed, resource utilization, and minimizing disruptions in the workflow. Such training could lead to the development of more effective workflows. The workflow updater requires global information to function effectively, which can become problematic as the context length increases. This limitation could be addressed by employing a rig or a hierarchical approach to more precisely identify errors or areas lacking efficiency, thereby facilitating more targeted updates and improvements within the workflow.
@@
  • Let $P_{A}(v)$ and $P_{B}(v)$ denote the probability that subtasks $v$ is successfully completed in Workflow A and Workflow B, respectively.
  • For each subtasks $v$, let $D_{A}(v)$ and $D_{B}(v)$ be the sets of immediate predecessors of $v$ in Workflow A and Workflow B, respectively.
Success Probability of a subtasks: In Workflow A, the success probability of subtasks $v$ is given by:
$P_{A}(v)=(1-p_{f})\times\prod_{i\in D_{A}(v)}P_{A}(i).$ (1)
@@
$P_{B}(v)=(1-p_{f})\times\prod_{i\in D_{B}(v)}P_{B}(i).$ (2)
Base Case: Since the subtasks $v$ with no dependencies (i.e., $D_{A}(v)=D_{B}(v)=\emptyset$) have the same success probability in both workflows:
$P_{A}(v)=P_{B}(v)=1-p_{f}.$
@@
Comparison for Subtasks $v^{*}$: Subtasks $v^{*}$ has an additional dependency $d$ in Workflow B. Therefore:
$D_{B}(v^{*})=D_{A}(v^{*})\cup\{d\}.$
@@
$P_{A}(v^{*})$ $=(1-p_{f})\times\prod_{i\in D_{A}(v^{*})}P_{A}(i),$
$P_{B}(v^{*})$ $=(1-p_{f})\times\prod_{i\in D_{B}(v^{*})}P_{B}(i)=(1-p_{f})\times P_{B}(d)\times\prod_{i\in D_{A}(v^{*})}P_{B}(i).$
Since $D_{A}(v^{*})=D_{B}(v^{*})\setminus\{d\}$, and $P_{A}(i)=P_{B}(i)$ for all $i\neq v^{*}$ (because their dependencies are the same), it follows that:
$P_{B}(v^{*})=P_{A}(v^{*})\times P_{B}(d).$
@@
$P_{B}(v^{*})=P_{A}(v^{*})\times P_{A}(d)<P_{A}(v^{*}).$
Success Probabilities for Other Subtasks: For all subtasks $v\neq v^{*}$, $D_{A}(v)=D_{B}(v)$, so:
$P_{A}(v)=P_{B}(v).$
@@
$E[S_{B}]$ $=\sum_{v\neq v^{*}}P_{B}(v)+P_{B}(v^{*})$
$=\sum_{v\neq v^{*}}P_{A}(v)+P_{B}(v^{*})$
$=\left(\sum_{v\in\mathcal{T}}P_{A}(v)-P_{A}(v^{*})\right)+P_{B}(v^{*})$
$=E[S_{A}]-\left(P_{A}(v^{*})-P_{B}(v^{*})\right).$
Since $P_{B}(v^{*})<P_{A}(v^{*})$, the difference $\Delta P=P_{A}(v^{*})-P_{B}(v^{*})>0$. Thus,